Consider the situation of the previous problem. If the mirror reflects only 64% of the light energy falling on it, what will be the ratio of the maximum to the minimum intensity in the interference pattern observed on the screen?

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. Given that, the mirror reflects 64% of energy (intensity) of the light. So, l base 1/l base 2 = 0.64 = 16/25 ⇒ r base 1/r base 2 = 4/5 So, l base max/l base min = (r base 1 + r base 2)^2/( r base 1 - r base 2)^2 = 81: 1.